Journalist Report – March 3th

Sol 19 – Under the Spotlight

“All down the way the pursued and the pursuing, the dream and the dreamers”

– Chapter 19 of The Martian Chronicles by Ray Bradbury

After many Sols spent in the Martian storms, the conditions were ideal for the second photogrammetry EVA at Candor Chasma! The BBC film crew followed them for the first part of the EVA to North Ridge, before returning to the Hab to interview some of the crewmembers. The outing was a success: with only notes taken from the 2D map, Jérémy and Alexandre managed to find all 10 checkpoints in time. They came back to the Hab in time for a delicious lunch: a lentil Dahl prepared by Adrien, and a bowlful of cherry tomatoes from the GreenHab! I also took the time to prepare a batch of bread this morning, while Alice was HabCom and the others were out on EVA. I noticed how calming it was for me to focus on this task while letting my thoughts wander off for a little while, and talking to Alice when communication with the EVA crew was cut off. Moments shared with only one or two other crewmembers are rare, as we are almost always together during meals and after dinner, our only breaks during the day. Only late at night when most of the crew has gone off to sleep, or during moments like the one shared with Alice this morning, can we enjoy one-on-one conversations. I would say it is one of the things I miss the most about Earth!

In the afternoon, I visited Alice in the Science Dome, and she explained to me how she analyzed her rock samples gathered during the last EVA: using a stack of sieves with finer and finer meshes, she can separate the different size of grains present in a sandstone, and weigh the different parts to get an idea of the rocks’ composition. In the GreenHab, Adrien harvested spinach and showed me how well the plants in the aquaponics system had grown! The root system is now four times as dense as in the beginning of the mission. Aquaponics could be an interesting way of envisioning agriculture in space: no dirt is required, the water circulates within a closed loop, and the systems allows crops to be “planted” much closer to each other, optimizing the available space.

Back at the Hab, Corentin had prepared pancakes, and Quentin was analyzing the data from this morning’s EVA, retracing the crew’s steps through the canyon to better compare it with the next EVA, using the 3D maps!

EVA Report – March 3th

Crew 275 EVA Report 03-03-2023

EVA # 15

Author: Jeremy Rabineau (Commander and EVA Leader)

Purpose of EVA: The purpose of the EVA was to perform the photogrammetry experiments with the 2D map.

First, as it was on the way to Candor Chasma, we changed the batteries and collected the data of MegaAres, Purple Air, and the Field Mill. We also reinstalled the LOAC.

We then parked at North Ridge and explored the place, so that we could be filmed there by the BBC.

We then went to Candor Chasma where the two subjects of the experiment tried to find some specific checkpoints, while the operator of the experiment onsite took time measurements and drew their trajectory on a map. In addition, the two subjects were equipped with a step counter.

Start time: 09:01

End time: 12:49


09:01 Airlock closed.

09:06 Exiting the airlock.

09:14 The EVA Crew leaves the Hab with the rovers (Opportunity and Perseverance).

09:16 Rovers parked next to the location of the atmospheric instruments. The crew arrives onsite 3 minutes later.

09:28 Battery and SD Card of Purple Air changed.

09:29 LOAC installed and measurement launched.

09:33 Battery of the Field Mill Changed and measurement launched.

09:36 Data retrieved from MegaAres and battery changed.

09:38 Leaving with the rovers to North Ridge.

09:47 Arrival at North Ridge. Walking around the area so that the BBC can shoot beautiful videos.

10:16 The BBC crew is leaving, while the EVA crew starts walking towards Candor Chasma.

10:45 Arrival at the entrance of Candor Chasma.

11:47 All 10 checkpoints have successfully been identified by the two subjects of the experiment. Walking back towards Gateway to Candor.

12:29 EVA crew back at the rovers and driving back towards the MDRS.

12:40 Both rovers parked at the MDRS.

12:43 EVA Crew back to the main airlock.

Destination: First we stopped where the MegaAres and the METEO station are located, i.e., between the Hab and Marble Ritual; then we stopped at North Ridge; finally we went to Candor Chasma.

Coordinates (use UTM WGS 84): N4251000 E518000, N4252000 E519000, and N4251000 E521500

Participants: EVA Leader: Jeremy Rabineau (Crew Commander & Subject), Alexandre Vinas (Crew Astronomer & Subject), Quentin Royer (Crew Engineer & Operator)

Hab Com: Alice Chapiron (Crew Scientist)

Road(s) and routes per MDRS Map: Cow Dung Road until the intersection with “Gateway to Candor”, we then walked close to North Ridge, and then we went eastwards until Candor Chasma for an estimated 2.5 km from the rovers. (See map attached)

Mode of travel: Driving and walking.

GreenHab Report – March 3th

GreenHab Officer: Adrien Tison

Environmental control: Heater

Average Temperature: 71.1°F

Average humidity: 25%

Hours of supplemental light: 0 hour

Daily water usage for crops: 8.45 gallons

Daily water usage for research and/or other purposes: 0 gallon

Water in the blue tank: 119.42 gallons

Time(s) of watering for crops: 9:45

Changes to crops: I added sticks to support cucumbers, indeed, it was growing fast lately and taking too much space in the tank.

Narrative: As usual I went to maintain and water crops and plants in the morning.

Everything is looking good, greens I’ve planted a few days ago are starting to pop out of the dirt, and the beans I’ve also planted are looking good!

The GreenHab may be the most satisfying place in the station!

Surprisingly enough, cucumbers adapted quickly to this new positioning and their leaves are now facing the right way.

Harvest: 43g of spinach

Support/supplies needed: N/A

Operations Report – March 3th

Crew 275 Operations Report 03Mar2023
SOL: 19

Name of person filing report: Quentin Royer
Non-nominal systems:
• Toilet
• Spacesuit n°9

Notes on non-nominal systems: None

Spirit rover used: No
• Hours: 209.9
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Opportunity rover used: Yes
• Hours: 113.7
• Beginning charge: 100%
• Ending charge: 73%
• Currently charging: Yes

Curiosity rover used: No
• Hours: 221.7
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Perseverance rover used: Yes
• Hours: 255.9
• Beginning charge: 100%
• Ending charge: 61%
• Currently charging: Yes

General notes on rovers: Nothing to report.

Summary of Hab operations: The Crew conducted routine operations. Journalists from the BBC filmed some operations.
• WATER USE: 13.90
• Water: 404.27 gallons
• Static tank pipe heater: On
• Static tank heater: On
• Toilet tank emptied: No

Summary of internet: Hughes Net was used during the day before the opening of COMMS for email purposes, and Starlink (MDRSCampus-Guest) is used during COMMS. MDRScampus is used for experimental purpose.

Summary of suits and radios: Nothing to report.

Summary of GreenHab operations: The Crew botanist conducted routine operations. An interview of the GreenHab Officer was performed by BBC journalists.
• WATER USE: 8.45 gallons
• Heater: On
• Supplemental light: No
• Harvest: 43 g of spinach

Summary of ScienceDome operations: The Crew Scientist conducted experiments on rock samples. Interviews of the Crew Scientist and of the Commander were performed by BBC journalists.
• Dual split: Heat / From 10 pm to 2 am

Summary of RAM operations: No operations were conducted in the RAM today. An interview of the Crew Engineer was performed by BBC journalists.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns, and requests to Mission Support: None

Sol Summary Report – March 3th

Crew 275 Sol Summary Report 03Mar2023

Sol: 19

Summary Title: Under the spotlight

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

Finally, a day when we wake up and don’t see the landscape covered with snow or mud! The flipside of this coin is that the temperature in the spacesuits got higher than usual and the EVA crew was very tired after their 4-hour EVA. We are no longer used to walk 10 000 steps a day, and even less in a spacesuit! An exhausting EVA, but a very fulfilling one, as all the objectives have been accomplished in perfect agreement with the planned timeline. The LOAC has been re-installed, while the batteries and the SD cards of all the other atmospheric instruments have been changed. All the checkpoints of the photogrammetry experiment have been identified by the subjects onsite at Candor Chasma, where the terrain was still slightly muddy.

The main event of the day was the presence of a journalist crew from BBC. They accompanied us during the first hour of our EVA. They then filmed the activities in the Upper Deck, the lunch, and finally they spent the whole afternoon interviewing us at our dedicated workplaces: the Science Dome, the GreenHab, etc. It was honestly a bit strange to see other people after three weeks just among us, but we are happy that they were very cooperative and respectful regarding the simulation. We are looking forward to seeing the output they can get out of their footage.

Look Ahead Plan:

Tomorrow (Sol 20), we would like to have our third EVA to Candor Chasma. The EVA crews will have to find specific checkpoints based on a 3D render. We would like to take the opportunity of this EVA to change the batteries of the atmospheric instruments, since they cannot last for more than two days, and we are not planning any EVA for Sunday.

Another media visit is planned for next Tuesday (Sol 23) with a crew from France Televisions. We will keep you updated as soon as we have more information regarding their planned arrival and departure.

Anomalies in work: None.

Weather: Cloudy, low -5°C/23°F, high 8°C/46°F.

Crew Physical Status: Optimal.

EVA: EVA # 15 was successful, both the main (photogrammetry) and secondary (atmospheric instruments) objectives have been fulfilled.

Reports to be filed: Sol Summary, EVA Report, EVA Request, Operations Report, Journalist Report, GreenHab Report, Astronomy Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site